what is the frequency of kinetic energy of a body in SHM if the time period of SHM is T

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
From the motion of eqn.
y=Acos(wt+φ) 
=>dy/dt=-Asin(wt+φ) 
dy/dt=v 
The time of speed as an element of time will be the era of the sine bend that we got on the RHS of the condition. 
the day and age of the sine bend would be dependant on sin(wt) which is equivalent to 2π/w numerically. 
What's more, Kinteic vitality is equivalent to ½ x mv(t)2 
Consequently it would be dependant on sin2(wt) 
Scientifically if day and age of sinx is T then sin2x is T/2 (From chart) 
So KE will have day and age π/w 
What's more, we realize that recurrence = 1/day and age 
Recurrence of speed as a component of time=w/2π 
Recurrence of active vitality as an element of time=w/π 
Thus, f of SHM=1/T 
T=2π/w 
f of K.E=2/T.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)